OpenStudy (anonymous):

c+(4-3c)-2=0 help me solve this plez

OpenStudy (anonymous):

c + 4 - 3c -2 = 0 c - 3c = -4 + 2 -2 c = -2 c = 1

OpenStudy (dwelsher):

1. decide what you are solving for 2writ problem out with out the ( ) which in this case is straight forward 3. move like variables to one side of the equation and constants to the other side 4solve